OpenStudy (anonymous):

If f(x) = 3x and g(x) = 2x-f find (f o g)(x)

OpenStudy (anonymous):

Tell me if I'm wrong but dont you just replace what is on the gx onto the x of f(x)

OpenStudy (anonymous):

You are right

hero (hero):

f o g(x) = f(g(x)) f(x) = 3x g(x) = 2x-f f(g(x) = 3(2x-f)

OpenStudy (anonymous):

Oh ok f o g mean f of g so if you have f(x) f o g is now f(gx)

OpenStudy (anonymous):

f(x) means f of x if you didn't know before

OpenStudy (anonymous):

A. = -16x B. = 6x-6 C. = 6x-18

hero (hero):

Romero, if you want, you can also explain what f of x means as well.

hero (hero):

If those are the answer choices, then I'm 100% certain that you wrote the original equation for g(x) incorrectly @Lib Libs

hero (hero):

f o g(x) = f(g(x)) f(x) = 3x g(x) = 2x-6 f(g(x) = 3(2x-6) = 6x - 18

hero (hero):

That's just my guess at what g(x) is, but you can clear it up by stating that it is

OpenStudy (anonymous):

you explain what f of x means

OpenStudy (anonymous):

Thank you!~

hero (hero):

Libs, can you verify the correct expression for g(x) ?

hero (hero):

Please post your problem again. I don't believe that g(x) = 2x-f

hero (hero):

What I posted above was a complete assumption

OpenStudy (anonymous):

If f(x) = 3x and g(x) = 2x-f find (f o g)(x)